### v1

@@ 3.2 Data Analysis
The analysis for RQ1 involved calculating descriptive statistics about the distribution of talks, speakers, and sponsors. Answering RQ2 requires natural language processing (NLP) techniques, as illustrated in Figure 1. We followed a deductive approach by mapping abstracts to established security frameworks, which describe different areas and aspects of defensive and offensive security. The two main decisions were which frameworks and NLP techniques to use.
@@
Choice of Framework The framework would ideally be: (i) comprehensive, (ii) widely used in practice; and (iii) freely available. The first criterion ensures we can meaningfully map more abstracts to the framework, which would not be possible something like a digital forensics standard. The second criterion makes our results easier to understand as some readers will be familiar with the categories from the framework. The final criterion enables others to replicate our work.
@@
For offensive frameworks, MITRE ATT&CK and Cyber Kill Chain framework were considered the most appropriate choices. The ATT&CK framework offers greater depth, satisfying the first criterion. For defensive frameworks, we considered the NIST Cybersecurity Framework, ISO 27001 and CIS Critical Security Controls. We excluded ISO 27001 due to the cost of access. The CIS standard was excluded because NIST CSF is broader, including governance. Although the choice of framework is important, mappings between frameworks are widely available [^4] For example, CIS Controls v8 mapped to NIST CSF is available here: https://www.cisecurity.org/insights/white-papers/cis-controls-v8-mapping-to-nist-csf.
@@
NLP Techniques For the first classification phase, we employed OpenAI’s gpt-3.5-turbo model to classify abstracts into offensive, defensive, or neutral via this prompt:
@@
“is this text cyber security offensive or defensive (just answer one word, namely: offensive, defensive or neutral)? $<$abstract$>$
Neutral talks, which make up a minority (17%) of the corpus, are not further classified.
@@
For the second classification, we referred to the official documentation of each standard to establish themes and sub-themes. For MITRE ATT&CK Framework, we used the 14 tactics as themes and the 196 techniques as sub-themes. For NIST Cybersecurity Framework, we used the 22 categories as themes and the 98 sub-categories as sub-themes. We estimated the similarity between the abstract and the corresponding (sub) theme using two NLP techniques, namely cosine similarity and a Large Language Model.
To calculate cosine similarity, we converted abstracts and each (sub-)theme to a bag of words. We removed stop words using the python nltk stopwords list and then used the nltk wordnet lemmatizer to group the different word forms together. We then compared the talk to each category from the framework using cosine similarity. To compute this, we normalized embeddings with the SentenceTransformer model ‘stsb-mpnet-base-v2’. This process was repeated for the sub-categories of the NIST Cybersecurity Framework and also the techniques from MITRE ATT&CK Framework.
@@
For the Large Language Model, we again used OpenAI’s gpt-3.5-turbo model using the following prompt:
@@
“What are the MITRE ATT&CK tactics that can be identified in the following text. Only return the tactics which are identified: $<$abstract$>$
@@
We used a similar prompt for the categories of NIST CSF. Each theme was individually submitted as a separate query to the model, which produced either a 0 or 1 as output for each talk. To calculate aggregate similarity to each theme, we took the average across all talks in a specific sub-sample.
3.3 Validation
The absence of an existing ground truth mapping for abstract talks to cyber security frameworks made validation difficult. Further, frameworks are high-level with blurry boundaries between categories. To address these issues, we first evaluated the first classification using a hand-labeled data set. Second, we conducted a comparative analysis of the agreement between the two natural language processing techniques. Third, we measured the similarity of offensive, neutral, and defensive talks against both frameworks.
First, we used a human-labeled data set comprising 50 abstracts to evaluate the performance of OpenAI’s gpt-3.5-turbo model at classifying abstracts as offensive, defensive, or neutral. The machine and human agreed on 98% of abstracts.
@@ 3.3 Validation
We then compared the topic classifications across NLP techniques. Since GPT outputs are either 0 or 1, we first apply a threshold to the cosine similarity scores, which are floats between 0 and 1. We set the threshold such that both GPT and cosine similarity would have the same number of 0s and 1s. Otherwise the scores could be perfectly aligned, but appear to disagree purely due to setting different thresholds. For the ATT&CK and NIST Cybersecurity framework categories, the two techniques agreed for 67% and 81% of the abstracts respectively. This ranged across individual categories, from a low of 52% up to 95% (see Table 3). The categories with high accuracy are partly driven by the class imbalance as so few defensive abstracts cover Maintenance (from NIST CSF) or Resource Development (from MITRE ATT&CK).
@@
Third, it might be expected that offensive or defensive talks display the highest similarity to the offensive or defensive frameworks, while neutral talks would exhibit low similarity to both frameworks. Table 2 shows that the cosine similarity scores display the expected ranking. However, this is not true for OpenAI’s gpt-3.5-turbo model, as offensive talks display higher similarity to both frameworks, and defensive talks display the lowest similarity to NIST CSF.
@@
Table 2: Average similarity of each category of talks to each framework using both NLP techniques.
Summary Overall, these three validation tests suggest our mapping captures some signal, albeit with considerable noise. Some noise is to be expected given the subjective nature of the task. The GPT method reliably classifies talks as offensive, defensive or neutral. However, the agreement between NLP techniques is disappointing.
@@
Going forward, we will report on aggregate results from OpenAI’s gpt-3.5-turbo because it is a state-of-the-art LLM. Aggregating the classifications helps mitigate random errors as these should cancel out. However, it does not mitigate systemic bias associated with GPT, which we discuss in Section 5. In the Appendix, we also display figures based on an ensemble model, which outputs an equal weighting of the GPT and cosine similarity techniques.
@@
Table 3: Fraction of abstracts per category for which the GPT and cosine similarity techniques agree.
3.4 Open Science & Ethics
We will release our corpus of titles, abstracts, speakers, and sponsors in an open source repository, as well as our analysis scripts. We hope this allows other researchers to extend this research project. We discuss these directions in Section 6.
Our main ethical concern is that we process and then share personal data (e.g. author names and affiliations). It would be impracticable to collect affirmative consent from speakers. We believe the benefits in terms of open science outweigh the potential privacy invasion, not least because presenting at conferences is a public act.
4 Results
@@
4.1 Conference Statistics
RSA and Black Hat had the most talks across the sample period (see Figure 2). Volatility in the size of Black Hat since 2019 is driven by the conference introducing/discontinuing sessions like Day Zero, CISO Summit, and Sponsored Workshops. The academic conferences display the highest rates of growth, as noted by Balzarotti balzarot2022 . USENIX Security doubled in size from 2019 to 2021. The industry conferences do not display the same growth trend. Insurance conferences had the fewest talks, with the absolute number remaining stable over time. This is perhaps surprising given the market grew by over 360% between 2015 and 2021 naic2022report .
Figure 2: Number of talks at each conference over time.
Figure 3: Top sponsors of academic conferences. The number on top of each bar is the mean sponsorship tier of that sponsor.
The academic conferences all had the longest abstracts in terms of word count, which can be considered an imperfect proxy for information content. Table 1 shows that IEEE S&P displays the highest average at 223 words. Abstracts at Black Hat and DEFCON had a similar word count to the other three academic conferences. These five conferences have a mean word count of between 170 and 190, which makes sense given original research is often presented at Black Hat and DEFCON. The conferences with the lowest word counts place less emphasis on research. There was less variance in the length of titles, both within and between conferences.
@@ 4.1 Conference Statistics
Figure 4: Speaker affiliation and sponsor correlation of Black Hat.
The academic conferences are sponsored by a mixture of public and private sources (see Figure 3). Large global technology companies, along with the National Science Foundation, support all four conferences in a similar proportion. InfoSec vendors tend to support specific conferences. Non-profits like the EFF, the FreeBSD Foundation, and DMTF only sponsor USENIX Security.
@@
The industry conferences are mainly sponsored by InfoSec vendors, with large technology companies sponsoring less frequently. This can be seen in Figure 19 in the Appendix. The analysis of sponsors of industry conferences is less comprehensive because of gaps in the data. The insurance conferences contain a wide range of sponsors (see Figure 19 in the Appendix). These include not just insurance companies, but also law firms, digital forensics providers, communications consultants, and data analytics firms. This is likely because these firms receive work from insurers woods2023lessons .
@@
Figure 5: Talks per individual for the top 30 most frequent speakers.
4.2 Speaker Statistics
This subsection examines the distribution of talks among specific individuals. Throughout we assume that each unique name corresponds to a single unique person. Figure 6 shows there is minimal overlap of speakers between industry, academic, and insurance conferences. For this reason, we proceed by analyzing each type of conference separately.
Figure 6: Number of unique speakers at each type of conference.
@@ 4.2 Speaker Statistics
Table 1 displays the Gini index for each conference. A score of 0 means all speakers gave the same number of talks, while a score of 1 means one individual gave all talks. Academic conferences are similarly equal to industry conferences when publications count as $\frac{1}{n}$ where $n$ is the number of co-authors/co-panelists, whereas academic conferences appear considerably less equal when each publications counts as $1$ (see Table 1). This corrects for academics who make small contributions to high $n$ papers—we all have just 24 hours in the day. Both metrics agree that NDSS is the most equal academic conference, while ACM CCS is the least.
Figure 7: Offensive, defensive, and neutral talks at each conference.
4.3 Content of Talks
We now explore the topics of abstracts. Figure 7 shows that defensive talks are more common than offensive and neutral talks at all conferences apart from DEFCON, which is explicitly set up for hackers to present their research. At DEFCON, 64% of the talks were offensive and just 15% (20%) of the talks were defensive (neutral). Across both cyber insurance conferences, just 1% of talks were classified as offensive, whereas 55% of insurance talks were neutral and 44% were defensive. The academic conference all accepted more defensive work than offensive, and some even accepted more neutral than offensive talks.
4.3.1 Defensive Talks
@@
Figure 8: Similarity score of defensive talks mapped to NIST Cybersecurity Framework over years
@@
Figure 9: Similarity score of DEFCON defensive talks mapped to NIST Cybersecurity Framework over years
@@
The identify function comprises the first six rows in Figure 10. Broadly speaking, Identify covers how organizations make security decisions. Both industry InfoSec and insurance talks address these topics more than the academic talks. The insurance conferences and Gartner display high similarity to Risk Assessment, which captures the likelihood and impact of different security incidents, and Governance, which cover the process of establishing a security policy. It is also notable that the research-oriented industry information security conferences (DEFCON and Black Hat) display less similarity to Identify. Notably, there is low similarity to the final two sections: Risk Management Strategy involves understanding the firm’s meta-reasoning like the assumptions and risk tolerance of the firm, meanwhile Supply Chain Risk Management involves assessing the risk of external firms.
Researchers should not dismiss the Identify function as ‘non-technical’ and outside the scope of computer security. Many of the Identify tasks can be solved using technical data. For example, a recent SoK on asset discovery highlights how Asset Management can be approached using network measurements vermeer2021sok . Similarly, Risk Assessment is analogous to cyber incident prediction, which has been studied as a machine learning problem based on technical indicators and published at top tier venues liu2015cloudy ; bilge2017riskteller .
Figure 10: Similarity score of defensive talks mapped to NIST Cybersecurity Framework.
@@
Academic talks display some similarity to Identity Management and Access Control. This include the evergreen research topic that is passwords morris1979password ; bonneau2012quest . The Awareness and Training category received more coverage in the industry talks relative to the academic talks. Academic research related to this may instead be presented at conferences focusing on Human-Computer Interaction like SOUPS and Chi garfinkel2014usable .
The remaining three categories had low similarity. Maintenance had the lowest similarity across all of the defensive categories. This finding supports the observation that researchers often overlook maintenance even though it is crucial to successful operations collier2020cybercrime . The Information Protection Processes and Procedures category consists of procedures to ensure availability like back-ups, recovery plans and so on. Industry talks more frequently cover these topics. Low similarity to Data Security was perplexing given this is so core to computer security. The first two sub-categories of Data Security, namely data at rest and in transit, are core academic topics. However, the other sub-categories are organizational procedures that tend not to be covered at academic conferences.
Figure 11: Cosine similarity score for the category within Protective Technology.
@@
The respond function comprises the next five rows, which are broadly the steps taken to investigate and contain an incident. Abstracts were most similar to Response Planning, which involves establishing incident response procedures before the incident occurs. The talks also displayed reasonable similarity to Mitigation, which involves actions to contain and resolve an incident, and also Improvements, which is about extracting lessons learned from incidents. This is perhaps surprising given the consensus that the security community struggles to extract lessons from incidents knake2021learning .
There was low similarity with Analysis category, which includes forensic investigations and vulnerability disclosure policies, especially at academic conferences. This could be because researchers present at specialist venues like Digital Investigation garfinkel2010digital . There was relatively low similarity to Communications, which involves coordinating with internal and external stakeholders like law enforcement, especially at academic conferences.

### v2

@@ 3.2. Data Analysis
The analysis for RQ1 involved calculating descriptive statistics about the distribution of talks, speakers, and sponsors. To address RQ2, we applied natural language processing (NLP) methods to classify abstracts according to established cybersecurity frameworks (Figure 1). This deductive approach mapped textual content to structured taxonomies of offensive and defensive security practices. Two key decisions guided this analysis: (i) the selection of frameworks and (ii) the choice of NLP techniques.
@@
Choice of Framework The chosen frameworks would ideally be: (i) comprehensive, (ii) widely recognized in practice, and (iii) freely accessible. The first criterion ensures us to meaningfully map more abstracts to the framework, which would not be possible with a digital forensics standard. The second criterion makes our results easier to understand, since some readers will always be familiar with the categories in the framework. The final criterion enables others to replicate our work.
@@
For offensive frameworks, we considered MITRE ATT&CK and Cyber Kill Chain framework. ATT&CK framework offers greater depth, satisfying the first criterion. For defensive frameworks, we evaluated NIST Cybersecurity Framework (CSF), ISO 27001 and CIS Critical Security Controls. We excluded ISO 27001 due to licensing costs. CIS standard was excluded because NIST CSF is broader, including governance. Although the choice of framework is important, mappings between frameworks are widely available [^5] For example, CIS Controls v8 mapped to NIST CSF is available here: https://www.cisecurity.org/insights/white-papers/cis-controls-v8-mapping-to-nist-csf (accessed 19 October 2025).
@@
NLP Techniques We implemented a two-stage classification process. For the first classification phase, we used OpenAI’s gpt-3.5-turbo model to classify abstracts as offensive, defensive, or neutral using the following minimal zero-shot prompt:
@@
“Is this text cybersecurity offensive or defensive (just answer one word, namely: offensive, defensive or neutral)? <<abstract>>
Minimal prompting reduced output variation and improved scalability, as more detailed prompts in pilot tests led to longer, less consistent, and more ambiguous responses. Neutral talks, which comprise a minority (17%) of the corpus, are not further classified, as they can not be clearly mapped to either frameworks (see Table 4). They are mentioned solely for completeness.
@@
For the second stage of the classification, we referred to the official documentation of each standard to establish themes and sub-themes. For MITRE ATT&CK Framework, we used the 14 tactics as themes and the 196 techniques as sub-themes. For NIST Cybersecurity Framework, we used the 22 categories as themes and the 98 sub-categories as sub-themes. We estimated the similarity between each abstract and the corresponding (sub) theme using two NLP techniques: cosine similarity and a Large Language Model (LLM).
To calculate cosine similarity, we preprocessed text by removing stop words (NLTK stopword list) and lemmatized terms using WordNetLemmatizer. Texts were represented as embeddings via SentenceTransformer (stsb-mpnet-base-v2) and compared to framework categories using cosine similarity. This process was applied to both NIST CSF and MITRE ATT&CK hierarchies.
@@
For the Large Language Model, we used OpenAI’s gpt-3.5-turbo model using the following zero-shot prompt:
@@
“What are the MITRE ATT&CK tactics that can be identified in the following text. Only return the tactics which are identified: <<abstract>>
@@
A similar prompt was used for the NIST CSF categories. Each theme was queried individually, producing binary (0/1) outputs per talk. To calculate aggregate similarity to each theme, we took the average across all talks in a specific sub-sample.
3.3. Validation
Validation was difficult because there is no existing ground truth mapping for abstract talks to cybersecurity frameworks. Further, frameworks are high-level with blurry boundaries between categories. We employed a three-pronged validation approach to address these issues, consisting of: (i) human annotation, (ii) cross-method agreement, and (iii) theoretical consistency.
For human annotation, we manually labeled 264 abstracts and compared these to GPT classification. The model achieved a $92.4\%$ concordance with the human annotators. Based on the selected sample ($n=264$) from the total population of $N=9,728$ abstracts, a $95\%$ confidence interval for true population concordance was determined at $\pm 3.16$ percentage points. This calculation incorporated the Finite Population Correction (FPC) to accurately reflect the known, finite size of the total abstract population.
@@ 3.3. Validation
For the cross-method agreement, we compared the topic classifications across NLP techniques. Because cosine similarity produced continuous values, we applied a threshold to align the number of positive (1) and negative (0) classifications. We set the threshold such that both GPT and cosine similarity would have the same number of 0s and 1s. Otherwise the scores could be perfectly aligned, but appear to disagree purely due to setting different thresholds. For the ATT&CK and NIST Cybersecurity framework categories, the two techniques agreed for 67% and 81% of the abstracts respectively. This ranged across individual categories, from a low of 52% up to 95% (see Table 5). The categories with high accuracy are partly driven by the class imbalance as so few defensive abstracts cover Maintenance (from NIST CSF) or Resource Development (from MITRE ATT&CK).
@@
For the theoretical consistency, we examined whether offensive and defensive talks displayed similarity patterns. It might be expected that offensive or defensive talks display the highest similarity to the offensive or defensive frameworks, while neutral talks would exhibit low similarity to both frameworks. Table 4 compares the performance of the two classification techniques in aligning with our labeling of abstracts as offensive, defensive, or neutral. Specifically, it presents the average similarity scores between each category of talk and the corresponding cybersecurity frameworks. The results show that the cosine similarity scores display the expected ranking: defensive talks exhibit the highest similarity to the NIST CSF (0.317), and offensive talks align most closely with the ATT&CK framework (0.452). However, this is not true for OpenAI’s gpt-3.5-turbo model, as offensive talks display higher similarity to both frameworks, and defensive talks display the lowest similarity to NIST CSF.
@@
Table 4. Average similarity between talk categories and cybersecurity frameworks, computed using GPT-based classification and cosine similarity techniques.
Summary These three validation tests indicates that our mapping captures meaningful signal, albeit with considerable noise - unsurprising given the subjective nature of the task. The GPT method reliably classifies talks as offensive, defensive or neutral. However, the agreement between NLP techniques is disappointing.
@@
Going forward, we will report on aggregate results from OpenAI’s gpt-3.5-turbo as it is a robust, widely used, and well-documented LLM variant. Aggregating the classifications helps mitigate random errors as these should cancel out. However, it does not mitigate systemic bias associated with GPT, which we discuss in Section 5. In the Appendix, we display supplementary figures based on an ensemble model, which outputs an equal weighting of the GPT and cosine similarity techniques.
@@
Table 5. Agreement between NLP techniques. Fraction of abstracts per category for which the GPT and cosine similarity classifications are consistent.
3.4. Open Science & Ethics
We release our corpus of titles, abstracts, speakers, and sponsors in an open source repository, as well as our analysis scripts (Walter et al., 2024). We hope this allows other researchers to extend this research project. We discuss these directions in Section 6.
Our primary ethical consideration is the use of personal data—such as author names and affiliations—which we process and analyze. As this information is already publicly available and tied to individuals’ voluntary participation in public events, we believe the privacy risks are minimal. While obtaining affirmative consent from all speakers would be impractical, we acknowledge the concern and have taken care to handle the data responsibly. We believe the value of this work for open science justifies the approach, though future collaborations with conference organizers could further strengthen ethical alignment.
4. Results
@@
4.1. Conference Statistics
RSA and Black Hat had the most talks across the sample period (see Figure 2). Volatility in the size of Black Hat since 2019 is driven by the conference introducing/discontinuing sessions like Day Zero, CISO Summit, and Sponsored Workshops. The academic conferences display the highest rates of growth, as noted by Balzarotti (Balzarotti, 2022). USENIX Security doubled in size from 2019 to 2021. The industry conferences do not display the same growth trend. Insurance conferences had the fewest talks, with the absolute number remaining stable over time. This is perhaps surprising given the market grew by over 360% between 2015 and 2021 (National Association of Insurance Commissioners Staff, 2021).
Figure 2. Number of talks presented at each conference from 2014 to 2022.
Figure 3. Top sponsors of academic conferences aggregated over time (2014–2022). Numbers above the bars indicate sponsorship tier of each sponsor.
The academic conferences all had the longest abstracts in terms of word count. This should be considered an imperfect proxy for information content, especially given many conferences have a word limit. Table 1 shows that IEEE S&P displays the highest average at 223 words. Abstracts at Black Hat and DEFCON had a similar word count to ACM CCS, NDSS, and USENIX Security. These five conferences have a mean word count of between 170 and 190, which makes sense given original research is often presented at Black Hat and DEFCON. The conferences (RSA, NetDiligence, and Advisen) with the lowest word counts place less emphasis on research. There was less variance in the length of titles, both within and between conferences.
@@ 4.1. Conference Statistics
Figure 4. Relationship between speaker slots and sponsor slots at the Black Hat conference (2014–2022). Each point represents a company, with the x-axis showing speaker slots and the y-axis showing sponsor slots.
The academic conferences are sponsored by a mixture of public and private sources (see Figure 3). Large global technology companies, along with the National Science Foundation, support all four conferences in a similar proportion. Information security (InfoSec) vendors tend to support specific conferences. Non-profits like the EFF, the FreeBSD Foundation, and DMTF only sponsor USENIX Security.
@@
The industry conferences are mainly sponsored by InfoSec vendors, with large technology companies sponsoring less frequently. This can be seen in Figure 19 in the Appendix. The analysis of sponsors of industry conferences is less comprehensive because of gaps in the data. The insurance conferences contain a wide range of sponsors (see Figure 19 in the Appendix). These include not just insurance companies, but also law firms, digital forensics providers, communications consultants, and data analytics firms. This is likely because these firms receive work from insurers (Woods et al., 2023).
@@
Figure 5. Number of talks per individual for the 30 most frequent speakers from 2014 to 2022 across industry, academic, and insurance conferences.
4.2. Speaker Statistics
This subsection examines the distribution of abstracts among specific individuals. At industry conferences, the individual typically presents at the conference, sometimes joint with others. However, just one author typically presents at academic conferences. We consider each type of talk separately to account for this, using “author” as the terms for individuals associated with academic talks. Throughout we assume that each unique name corresponds to a single unique person. Figure 6 shows there is minimal overlap of speakers between industry, academic, and insurance conferences. For this reason, we proceed by analyzing each type of conference separately.
Figure 6. Overlap of speakers across conference types. The Venn diagram shows intersections among industry, InfoSec, and insurance venues during 2014–2022.
@@ 4.2. Speaker Statistics
Table 3 displays the Gini index for each conference. A score of 0 means all speakers gave the same number of talks, while a score of 1 means one individual gave all talks. Academic conferences are similarly equal to industry conferences when publications count as $\frac{1}{n}$ where $n$ is the number of co-authors/co-panelists, whereas academic conferences appear considerably less equal when each publications counts as $1$ (see Table 3). This corrects for academics who make small contributions to high $n$ papers. Both metrics agree that NDSS is the most equal academic conference, while ACM CCS is the least.
Figure 7. Distribution of offensive, defensive, and neutral talks at each conference from 2014 to 2022.
4.3. Content of Talks
We now explore the topics of abstracts. Figure 7 shows that defensive talks are more common than offensive and neutral talks at all conferences apart from DEFCON, which is explicitly set up for hackers to present their research. At DEFCON, 64% of the talks were offensive and just 15% (20%) of the talks were defensive (neutral). Across both cyber insurance conferences, just 1% of talks were classified as offensive, whereas 55% of insurance talks were neutral and 44% were defensive. The academic conferences all accepted more defensive work than offensive, and some even accepted more neutral than offensive talks.
4.3.1. Defensive Talks
@@
Figure 8. Similarity score of defensive talks mapped to the NIST Cybersecurity Framework over time (2014–2022).
@@
Figure 9. Similarity score of DEFCON defensive talks mapped to the NIST Cybersecurity Framework over time (2014–2022).
@@
The identify function comprises the first six rows in Figure 10. Broadly speaking, Identify covers how organizations make security decisions. Both industry InfoSec and insurance talks address these topics more than the academic talks. The insurance conferences and Gartner display high similarity to Risk Assessment, which captures the likelihood and impact of different security incidents, and Governance, which cover the process of establishing a security policy. It is also notable that the research-oriented industry information security conferences (DEFCON and Black Hat) display less similarity to Identify. Notably, there is low similarity to the final two sections: Risk Management Strategy involves understanding the firm’s meta-reasoning like the assumptions and risk tolerance of the firm, meanwhile Supply Chain Risk Management involves assessing the risk of external firms. Academics study this area when they conduct threat modeling (Kaur et al., 2025; Usman and Zappala, 2025).
Researchers should not dismiss the Identify function as ‘non-technical’ and outside the scope of computer security. Many of the Identify tasks can be solved using technical data. For example, a recent SoK on asset discovery highlights how Asset Management can be approached using network measurements (Vermeer et al., 2021). Similarly, Risk Assessment is analogous to cyber incident prediction, which has been studied as a machine learning problem based on technical indicators and published at top-tier venues (Liu et al., 2015; Bilge et al., 2017).
Figure 10. Similarity score of defensive talks mapped to the NIST Cybersecurity Framework over time (2014–2022).
@@
Academic talks display some similarity to Identity Management and Access Control. This include the evergreen research topic that is passwords (Morris and Thompson, 1979; Bonneau et al., 2012). The Awareness and Training category received more coverage in the industry talks relative to the academic talks (Ho et al., 2025). Academic research related to this may instead be presented at conferences focusing on Human-Computer Interaction like SOUPS and Chi (Garfinkel and Lipford, 2014).
The remaining three categories had low similarity. Maintenance had the lowest similarity across all of the defensive categories. This finding supports the observation that researchers often overlook maintenance even though it is crucial to successful operations (Collier et al., 2020). The Information Protection Processes and Procedures category consists of procedures to ensure availability like back-ups, recovery plans and so on. Industry talks more frequently cover these topics. Low similarity to Data Security was perplexing given this is so core to computer security. The first two sub-categories of Data Security, namely data at rest and in transit, are core academic topics. However, the other sub-categories are organizational procedures that tend not to be covered at academic conferences.
Figure 11. Cosine similarity score for talks within the Protective Technology function of the NIST Cybersecurity Framework from 2014 to 2022.
@@
The respond function comprises the next five rows, which are broadly the steps taken to investigate and contain an incident. This does not capture coordinated response, such as botnet take downs (Vu et al., 2025). Abstracts were most similar to Response Planning, which involves establishing incident response procedures before the incident occurs. The talks also displayed reasonable similarity to Mitigation, which involves actions to contain and resolve an incident, and also Improvements, which is about extracting lessons learned from incidents. This is perhaps surprising given the consensus that the security community struggles to extract lessons from incidents (Knake et al., 2021).
There was low similarity with the Analysis category, which includes forensic investigations and vulnerability disclosure policies, especially at academic conferences. This could be because researchers present at specialist venues like Digital Investigation (Garfinkel, 2010). There was relatively low similarity to Communications, which involves coordinating with internal and external stakeholders like law enforcement, especially at academic conferences.